OpenStudy (anonymous):

this is embarrassing but how do you find the value of tan67.5˚ in surd form? Using the tan2θ identity and without a calculator?

OpenStudy (anonymous):

@TuringTest any idea?

OpenStudy (anonymous):

actually it's not embarrasing i don't know what surd form is lol xD

OpenStudy (anonymous):

uh its like, square roots and no decimals? i think you call them radicals

OpenStudy (turingtest):

it's not obvious to me how to do it immediately, but this will surely help http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf

OpenStudy (anonymous):

i have that >< i've filled in the tangent identity but i can't find a link to tan67.5˚

OpenStudy (turingtest):

well considering we are supposed to use the \(\tan2\theta\) identity, if we treat \(\theta\) as 67.5 deg, then \(2\theta=135^{\circ}\), which is a special value on the unit circle i would prefer to turn this into radians for simplicity

OpenStudy (anonymous):

ahhhh damn that was thankyu

OpenStudy (turingtest):

welcome :)